Perform the following base conversions using subtraction or division-remainder

3104 base 10 to base 9

4401 base to to base 9

1518 base 10 to base 7
Answer #1
1)
Divide 3104 successively by 9 until the quotient is 0
3104/9 = 344, remainder is 8
344/9 = 38, remainder is 2
38/9 = 4, remainder is 2
4/9 = 0, remainder is 4
Read remainders from the bottom to top as 4228
Answer = 4228

2)
Divide 4401 successively by 9 until the quotient is 0
4401/9 = 489, remainder is 0
489/9 = 54, remainder is 3
54/9 = 6, remainder is 0
6/9 = 0, remainder is 6
Read remainders from the bottom to top as 6030
Answer = 6030

3)
Divide 1518 successively by 7 until the quotient is 0
1518/7 = 216, remainder is 6
216/7 = 30, remainder is 6
30/7 = 4, remainder is 2
4/7 = 0, remainder is 4
Read remainders from the bottom to top as 4266
Answer = 4266